OpenStudy (anonymous):

sir in a Log how can we add the the value given below? 0.0899-bar 1.8435-bar 3.8751-3.1065=? plz comlete method

OpenStudy (anonymous):

if bar is with 1 then it means -1 and if bar is with 3 then it means -3

OpenStudy (loser66):

so, it is 0.0899-(-1.8435)-(-3.8751)-3.1065= 0.0899+1.8435+ 3.8751-3.1065 right?

OpenStudy (anonymous):

yes right.....

OpenStudy (loser66):

and do what? all = 2.702 how does it relate to log?

OpenStudy (anonymous):

in first value 0 is characteristics a and 0899 is mentisa and so on in all value

OpenStudy (anonymous):

in my key the answer is -2.9518

OpenStudy (loser66):

I am sorry. I am helpless here. @campbell_st

OpenStudy (loser66):

@surjithayer

OpenStudy (campbell_st):

quick question, you said logs so are the questions log(0.0899) - log(1.8435)..?

OpenStudy (anonymous):

no i want to ask simple is [log(0.7214)+log(20.37)-log(60.8)] its answer is -1.3833 i want to ask how?

OpenStudy (campbell_st):

well I used base 10 logs and got -0.616736

OpenStudy (loser66):

me too, XD

OpenStudy (campbell_st):

if you use a different base, you'll have a different answer. as an example base e logs gives -1.42009

OpenStudy (anonymous):

from base 10 .... Log(0.7214)=-1.8582 + log(20.37)=1.3089 - log(60.8)=1.7839 ?

OpenStudy (anonymous):

now........ -1.8582+1.3089-1.7839=-1.3833 the answer is confirm

OpenStudy (campbell_st):

well try it this way, using log laws \[\log_{10}(\frac{0.7214 \times 20.37}{60.8}) = \] wow for log(0.7214) I got -0.141824 log(20.37) I got 1.30899 and log(60.8) i got 1.7839 so there seems to be some difference.... unless there is a typo in the question. hope this helps